OpenStudy (anonymous):

9.02] What are the solutions of 6x2 − x − 15 = 0? x = 5 over 6, x = −3 x = −5 over 6, x = 3 x = 3 over 2, x = −5 over 3 x = −3 over 2, x =5 over 3

OpenStudy (anonymous):

Do you know how to use the quadratic formula?

OpenStudy (anonymous):

Yeah, but I don't understand it, sorry about that!

OpenStudy (anonymous):

Do you mind writing out the quadratic equation? I'll show you how to use it to solve this equation.

OpenStudy (anonymous):

You still there?

OpenStudy (anonymous):

Yep, I'm here! Ok, I will, wait a sec! :D

OpenStudy (anonymous):

Just google; quadratic formula.

OpenStudy (anonymous):

y = ax2 + bx + c

OpenStudy (anonymous):

Um, that is not really correct. look at the attachment.

OpenStudy (anonymous):

Can you view the attachment? If so, can you see the letters a, b, c? They are the ones that will get replaced with numbers from your equation: 6x2 − x − 15 = 0

OpenStudy (anonymous):

Yep, I saw it, but the lesson put this.... A quadratic equation is an equation of the form y = ax2 + bx + c, where “a”, “b” and “c” are real numbers

OpenStudy (anonymous):

Alright, then call this the quadratic formula.

OpenStudy (anonymous):

Now replace the a with 6, the b with -1, and c with -15. Do you know where i got these numbers?

OpenStudy (anonymous):

from the equation in the question.

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Can you draw that out? with the letters replaced with the numbers that I posted above?

OpenStudy (anonymous):

a with 6, the b with -1, and c with -15.

OpenStudy (anonymous):